14. Find the sum of the odd numbers between 0 and 50.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
The odd numbers between 0 and 50 are 1, 3, 5, 7, 9 … 49. Therefore, we can see that these odd numbers are in the form of A.P. Hence, First term, a = 1 Common difference, d = 2 Last term, l = 49 By the formula of last term, we know, l = a+(n−1) d 49 = 1+(n−1)2 48 = 2(n − 1) n − 1 = 24 n = 25 = Number of terms By the formula of sum of nth term, we know, Sn = n/2(a +l) S25 = 25/2 (1+49) = 25(50)/2 =(25)(25) = 625